OpenStudy (anonymous):

6x2/3 namelnemo answer

OpenStudy (mynameisnemo):

jesiljdlgjdlfjhldfjg just medal that lol i already did on the other

OpenStudy (anonymous):

i cant see other i close it

OpenStudy (mynameisnemo):

6/2 times 3/1 common deno 6/2 times 6/2 36/2 18

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

what is another way to look at this problem 3/4x16=12 A4/3x16=12 B3/4x16=12 C4x16/3=12 D3x16/4=12

OpenStudy (anonymous):

wich one a b c or d

OpenStudy (mynameisnemo):

give me a fricken sec

OpenStudy (anonymous):

ok u just got sassy

OpenStudy (mynameisnemo):

whats the difference between C and the original?

OpenStudy (mynameisnemo):

I meant B not C

OpenStudy (anonymous):

i dunna know so thats not it

OpenStudy (mynameisnemo):

well theyre the same thing

OpenStudy (mynameisnemo):

The answer is D though

OpenStudy (anonymous):

yea but it has to meen the same but look different

OpenStudy (mynameisnemo):

D!

OpenStudy (anonymous):

oh ok didnt see that ok

OpenStudy (anonymous):

bye see ya later i think

OpenStudy (mynameisnemo):

bye